OpenStudy (anonymous):

Medal! Find a solution to the problem -2x+5y=8 2x+2y=6

OpenStudy (hoblos):

add the 2 equations what will you get ?

OpenStudy (anonymous):

7y=14

OpenStudy (anonymous):

??

OpenStudy (hoblos):

yeah so y=?

OpenStudy (anonymous):

The choices are (-2,1) (2,1) (1,2) (0,3) though?

OpenStudy (hoblos):

i want give the final answer, i'm trying to guide you into how to solve it and you'll find the answer by yourself, we still have 2 easy steps 7y=14 solve for first then replace the value of y in the first equation and solve for x tell me what you get to check it for you

OpenStudy (anonymous):

-2x+5(2)=8 2x+2(2)=6 right?

OpenStudy (hoblos):

right now choose one of the two equations and solve for x

OpenStudy (anonymous):

2x+2(2)=6 2x+4=6/2 x+2=3 is that right so far?

OpenStudy (hoblos):

yeah so x=?

OpenStudy (anonymous):

x=1?

OpenStudy (hoblos):

yes good job

OpenStudy (anonymous):

so (1,2) is the answer?

OpenStudy (hoblos):

yes that's it

OpenStudy (anonymous):

thanks!!

OpenStudy (hoblos):

any time :)